7(3x+ 4) = 8(2x + 5) + 13
Answer by THANApHD(104)

7(3x+ 4) = 8(2x + 5) + 13
21x + 28 = 16x + 40 +13
21x - 16x = 53 - 28
5x = 25
x = 5